# An Efficient Sign Language Translation Using Spatial Configuration and Motion Dynamics with LLMs

### v1

@@ 3.4 Sign Adapter
In the previous sections, we extracted two distinct visual features: the spatial features $Z_{s}$ and the motion features $Z_{m}$. These features differ in both their dimensions and representation, as depicted in Figure 2. To effectively integrate these features, we introduce an additional module called Sign Adaptor (SA). As shown in Figure 3, SA includes linear projection layers, a 1D CNN, and a Multi-Layer Perceptron (MLP). These components work together to integrate the spatial and motion features into a unified sign representation, denoted as $Z_{sm}$. First, the spatial and motion features are passed through linear projection layers to transform them into features with matching dimensions. Next, the 1D CNN is applied for short-term modeling of the combined features. Finally, a cross-modal MLP Liu et al. (2024a) is employed to bridge the visual and textual modalities. The resulting outputs are represented as $Z_{sm}\in\mathbb{R}^{M\times d^{\prime}}$, where $M$ represents the reduced number of frames after convolution, and $d^{\prime}$ is the dimension aligned with that of the LLM. Although SA aids in bridging the modality gap between visual and textual features during training under the SLT supervision, the gap still persists. To tackle this issue, we introduce VT-Align, which will be detailed in the next section.
@@ 3.5 Visual-Text Alignment
VT-Align is a warm-up and go process designed to provide the SA module with well-initialized weights before the SLT supervision begins. This initial alignment is crucial, as it helps the model more effectively bridge the modality gap during training. To achieve this alignment, we use a widely-used softmax-based contrastive learning approach Radford et al. (2021); Jia et al. (2021).
Specifically, given a mini-batch =$\mathcal{B}=\{(S_{1},Y_{1}),(S_{2},Y_{2}),...\}$ of sign-text pairs, the contrastive learning objective encourages the embeddings of matching pairs $(S_{i},Y_{i})$ to align closely while pushing apart the embeddings of mismatched pairs $(S_{i},Y_{j\neq i})$. Text features $Z_{t}$ are extracted from the target translation $Y_{i}$ using the LLM’s embedding layer $E_{llm}(\cdot)$. Note that only the SA module $f_{sa}(\cdot)$ is updated during this process, while $E_{llm}(\cdot)$ remains fixed to preserve the LLM’s language capabilities. The VT-Align loss function $L_{vt}$ is represented as follows:
@@ 3.6 Training Details
Our framework is optimized in two stages: an initial warm-up phase followed by training with the SLT supervision. In the warm-up phase, we begin by training the SA module using VT-Align for a designated number of steps (e.g., 4K steps). After completing the warm-up phase, we proceed to a joint training of both SA and the LLM. For fine-tuning the LLM, we utilize LoRA Hu et al. (2021), a lightweight and efficient method specifically designed for this purpose. Overall, our method is trained with a combined loss function as:
$\mathcal{L}_{\textit{SpaMo}}=\mathcal{L}_{ce}+\lambda\mathcal{L}_{vt},$ (2)
where $\lambda$ is a hyperparameter, and $\mathcal{L}_{ce}$ represents cross-entropy loss.
Setting Methods Vis. Ft. BLEU-1 BLEU-2 BLEU-3 BLEU-4 ROUGE
Gloss-based SLRT Camgoz et al. (2020) $\checkmark$ 46.61 33.73 26.19 21.32 -
STN-SLT Voskou et al. (2021) $\checkmark$ 48.61 35.97 28.37 23.65 -
STMC-T Zhou et al. (2021b) 46.98 36.09 28.70 23.65 46.65
BN-TIN-Transf.+SignBT Zhou et al. (2021a) 50.80 37.75 29.72 24.32 49.54
SimulSLT Yin et al. (2021) 48.23 35.59 28.04 23.14 49.23
PET Jin et al. (2022) $\checkmark$ 49.54 37.19 29.30 24.02 49.97
MMTLB Chen et al. (2022a) $\checkmark$ 53.97 41.75 33.84 28.39 52.65
TS-SLT Chen et al. (2022b) $\checkmark$ 54.90 42.43 34.46 28.95 53.48
SLTUNET Zhang et al. (2023b) $\checkmark$ 52.92 41.76 33.99 28.47 52.11
Weakly Gloss-free TSPNet Li et al. (2020b) $\checkmark$ 36.10 23.12 16.88 13.41 34.96
GASLT Yin et al. (2023) $\checkmark$ 39.07 26.74 21.86 15.74 39.86
ConSLT Fu et al. (2023) $\checkmark$ - - - 21.59 47.69
Gloss-free CSGCR Zhao et al. (2021) 36.71 25.40 18.86 15.18 38.85
GFSLT-VLP Zhou et al. (2023) $\checkmark$ 43.71 33.18 26.11 21.44 42.29
FLa-LLM Chen et al. (2024) $\checkmark$ 46.29 35.33 28.03 23.09 45.27
Sign2GPT Wong et al. (2024) $\checkmark$ 49.54 35.96 28.83 22.52 48.90
SignLLM Gong et al. (2024) $\checkmark$ 45.21 34.78 28.05 23.40 44.49
@@
Table 1: Performance comparison on the PHOENIX14T dataset. “Vis. Ft.” denotes to the visually fine-tuned on sign language datasets. The best results are highlighted in bold, and the second-best are underlined.
@@ 4.1 Implementation Details
For SE and ME, we use CLIP ViT-L/14 Radford et al. (2021) and VideoMAE-L/16 Tong et al. (2022), respectively. To extract the spatial features, the sign images are interpolated to multiple scales, such as $224\times 224$ and $448\times 448$. For each scale, larger images are split into sub-images of regular size ($224\times 224$) and processed individually. These features from the sub-images are then pooled and concatenated with features from the original representation. For the motion features, each clip consists of 16 frames aligned with the findings from Wilbur (2009), which suggests that this frame interval captures a single sign. We use an 8-frame gap between neighboring clips. We utilize FlanT5-XL-16bit Chung et al. (2024) as our LLM. During the warm-up phase with VT-Align, we use 4K steps on PHOENIX14T and 15K steps on How2Sign. Additional implementation details can be found in Appendix Section A.
4.2 Datasets and Evaluation Metrics
@@
We evaluated our method on two sign language datasets: PHOENIX14T Camgoz et al. (2018) and How2Sign Duarte et al. (2021). PHOENIX14T is a German Sign Language (DGS) dataset focused on weather forecasts, featuring a closed domain with a vocabulary of 3K words and an average of 116 frames. How2Sign is a large-scale American Sign Language (ASL) dataset that spans over a more open instructional domain, containing a vocabulary of 16K words and an average of 173 frames. Detailed statistics for both datasets are provided in Appendix Section C.
@@
We report BLEU via SacreBLEU Papineni et al. (2002); Post (2018)222nrefs:1|case:mixed|eff:no|tok:13a|smooth:exp|version:2.2.1 and ROUGE-L Lin and Och (2004). BLEU-n assesses translation precision by evaluating n-grams. ROUGE-L measures text similarity by calculating the F1 score based on the longest common subsequences between predicted and reference texts. We also report BLEURT Sellam et al. (2020) from the BLEURT-20 checkpoint [^3] https://huggingface.co/lucadiliello/BLEURT-20, which has been shown to correlate well with human judgments.
@@ 4.3 Comparison with State-of-the-Art
A Taxonomy of SLT.
In Section 2, we explore gloss-free methods, including those that incorporate gloss-supervised visual encoders. Although these approaches have traditionally been categorized as gloss-free, we argue that they should more accurately be described as weakly gloss-free due to their dependence on gloss-annotated data. This classification is detailed in Table 3.6. Specifically, methods such as TSPNet Li et al. (2020b), GASLT Yin et al. (2023), ConSLT Fu et al. (2023), GloFE-VN Lin et al. (2023), and OpenSLT Tarrés et al. (2023) rely on sign features extracted by visual encoders trained on continuous or isolated sign language recognition (SLR) datasets.
Results on PHOENIX14T.
We compare our method with both gloss-based and gloss-free methods on PHOENIX14T. As shown in Table 3.6, most previous methods rely on the domain-specific fine-tuning of their visual encoders. By contrast, our method demonstrates consistent improvements across all reported metrics without such fine-tuning, except for ROUGE, where it achieves the second-best result. Notably, the improvement on BLEU-4 is particularly significant, with a margin of 0.92, representing a 3.93% increase over SignLLM Gong et al. (2024). These results demonstrate that conveying the key components of sign language to LLMs is crucial for enhancing translation performance and adopting LLMs in SLT without expensive training or complex training steps.
@@
We also evaluated our method on How2Sign, which poses greater challenges than PHOENIX14T due to its more open-domain nature, longer sign video lengths, and larger vocabulary. The results are presented in Table 3.6. Our method outperforms previous methods across all reported metrics. Specifically, we achieve a 0.45 margin in BLEU-4, representing a 4.66% improvement over Fla-LLM Chen et al. (2024). Notably, we see a performance gain in BLEURT, reaching 2.93, which is 7.46% higher than SSVP-SLT Rust et al. (2024) under the non-scaled dataset setting. These results suggest that our method is more robust for open-domain scenarios with general topics and longer video lengths, even without the domain-specific fine-tuning.
@@
Table 3: Ablation study of main component.
Models Params B1 B2 B3 B4 RG
W/o LLM 60.5M 24.93 12.76 8.45 6.35 18.96
mT5-Large 1.2B 32.31 19.23 13.21 9.87 26.32
Flan-T5-Large 0.8B 47.63 34.75 27.04 22.02 43.66
@@
Table 4: Ablation study for impact of LLM.
@@
Figure 4: Ablation study for SE and ME. $S^{2}$ represents Scaling on Scales, and $n$ denotes the gap between neighboring clips. Note that the results presented do not include VT-Align.
@@
To assess the quality of sign representations, following Ye et al. (2023), we employ Kernel Density Estimation (KDE) to estimate the probability density functions of embeddings from GFSLT-VLP and our method on PHOENIX14T. Note that we reproduced GFSLT-VLP using the official code [^4] https://github.com/zhoubenjia/GFSLT-VLP. As shown in Table 5, our method produces more compact and confident representations than GFSLT-VLP. More details can be found in Appendix Section A
@@ 4.4 Ablation Study
To further evaluate our method, we conducted extensive ablation experiments on PHOENIX14T, the most widely used sign language dataset. Additional results can be found in Appendix Section B.
@@
We begin by evaluating the effect of the key components in our framework, specifically SE (spatial features), ME (motion features), and VT-align. As shown in Table 4.3, using both spatial and motion features alone yields results comparable to Sign2GPT in terms of BLEU-4 score (22.52 vs. 22.26). However, when VT-align is incorporated with these features, it achieves the best overall performance, demonstrating the importance of each component in SpaMo.
@@
Next, we examine the effect of different LLMs by replacing the model with various LLM types, as shown in Table 4.3. We compare four baselines, each with a different model size: our method without pre-trained weights, mT5-Large, Flan-T5-Large, and Flan-T5-XL. Of these, Flan-T5-XL demonstrates the best performance.
@@
Finally, we evaluate the effect of $S^{2}$ and the gap between neighboring clips on SE and ME, respectively. As shown in Figure 4, $S^{2}$ substantially improves translation performance, highlighting its effectiveness to help SE capture more detailed spatial details. Additionally, our analysis reveals that an 8-frame gap between neighboring clips produces the best results, indicating that this specific gap optimally facilitates ME in extracting dynamic motion information.
@@
Figure 5: The t-SNE visualization of sign features. Different colors represent sign features with distinct semantics, while gray points are other categories not listed.
@@ 4.5 Qualitative Analysis
Table 6 presents two example translations on PHOENIX14T, comparing our method with GFSLT-VLP, the only other publicly available baseline. In the first example (top), our method provides an accurate translation, whereas GFSLT-VLP fails to capture the correct semantic meaning. In the second example (bottom), our method again produces a precise translation, while GFSLT-VLP introduces errors, resulting in incorrect information. These examples demonstrate the superior accuracy of our method in generating reliable translations. Please refer to Appendix Section D for more translation examples.
@@
We performed an additional analysis to explore how the LLM interprets the sign videos. Inspired from the reverse engineering Ju et al. (2023), we first compute the Euclidean distance between the sign feature $Z_{sm}$ and the LLM’s embedding table $E_{llm}\in\mathbb{R}^{V\times d^{\prime}}$, where $V$ represents the vocabulary size. Each sign feature is then mapped to the word associated with the shortest distance in this space. This process can be expressed as $\text{dist}(Z_{sm},E_{llm})\leq\Delta$, where $\text{dist}(\cdot)$ denotes the Euclidean distance function, and $\Delta$ represents the shortest distance to $E_{llm}$ across all sign features.
Figure 5 shows the t-SNE visualization of each sign feature mapped to the corresponding vocabulary. We observed that certain visual features align closely with specific words, which likely represent the semantic concepts the LLM associates with these features. In other words, these words represent the LLM’s interpretation or labeling of the visual content. We refer to these mapped words as “visual tokens”. We further compare these visual tokens with the ground-truth glosses as shown in Table 7. To ensure a clearer and more accurate semantic comparison, repetitive words were removed from the visual tokens. Surprisingly, the LLM’s interpretation of the sign videos is similar to the glosses, though not perfectly aligned. This suggests that the LLM has learned to link particular video patterns with specific textual concepts, explaining why those words cluster near the visual features in the embedding space. Additionally, we found that the visual tokens capture words that are present in the translation but not in the glosses. This finding suggests that visual tokens may provide a more comprehensive representation than current glosses, potentially broadening their scope beyond what has been traditionally documented.

### v2

@@ 3.4 Sign Adapter
In the previous sections, we extracted two distinct visual features: the spatial features $Z_{s}$ and the motion features $Z_{m}$. These features differ in both their dimensions and representation, as depicted in Figure 2. To effectively integrate these features, we introduce an additional module called Sign Adaptor (SA). As shown in Figure 3, SA includes linear projection layers, a 1D CNN, and a Multi-Layer Perceptron (MLP). These components work together to integrate the spatial and motion features into a unified sign representation, denoted as $Z_{sm}$. First, the spatial and motion features are passed through linear projection layers to transform them into features with matching dimensions. Next, the 1D CNN is applied for short-term modeling of the combined features. Finally, a cross-modal MLP Liu et al. (2024a) is employed to bridge the visual and textual modalities. The resulting outputs are represented as $Z_{sm}\in\mathbb{R}^{M\times d^{\prime}}$, where $M$ represents the reduced number of frames after convolution, and $d^{\prime}$ is the dimension aligned with that of the LLM. Although SA aids in bridging the modality gap between visual and textual features during training under the SLT supervision, the gap remains. To tackle this issue, we introduce VT-Align, which will be detailed in the next section.
@@ 3.5 Visual-Text Alignment
VT-Align is a warm-up and go process designed to provide the SA module with well-initialized weights before the SLT supervision begins. This initial alignment is crucial, as it helps the model more effectively bridge the modality gap during training. To achieve this alignment, we employ a widely-used softmax-based contrastive learning approach Radford et al. (2021); Jia et al. (2021).
Specifically, given a mini-batch =$\mathcal{B}=\{(S_{1},Y_{1}),(S_{2},Y_{2}),...\}$ of sign-text pairs, the contrastive learning objective encourages the embeddings of matching pairs $(S_{i},Y_{i})$ to align closely while pushing apart the embeddings of mismatched pairs $(S_{i},Y_{j\neq i})$. Text features $Z_{t}$ are extracted from the target translation $Y_{i}$ using the LLM’s embedding layer $E_{llm}(\cdot)$. Note that only the SA module $f_{sa}(\cdot)$ is updated during this process, while $E_{llm}(\cdot)$ remains fixed to preserve the LLM’s language capabilities. The VT-Align loss function $\mathcal{L}_{vt}$ is represented as follows:
@@ 3.6 Training Details
Our framework is optimized in two phases: an initial warm-up phase followed by training with the SLT supervision. In the warm-up phase, we begin with a training phase the SA module using VT-Align for a designated number of steps (e.g., 4K steps). After completing the warm-up phase, we proceed to the joint training for both SA and the LLM. For fine-tuning the LLM, we utilize LoRA Hu et al. (2021), a lightweight and efficient method specifically designed for this purpose. Overall, our method is trained with a combined loss function:
$\mathcal{L}_{\textit{SpaMo}}=\mathcal{L}_{ce}+\mathcal{L}_{vt},$ (2)
where $\mathcal{L}_{ce}$ represents cross-entropy loss, and $\mathcal{L}_{vt}$ continuously manages the alignment.
PHOENIX14T CSL-Daily
Setting Methods Vis. Ft. B1 B2 B3 B4 RG B1 B2 B3 B4 RG
Gloss-based SLRT Camgoz et al. (2020) $\checkmark$ 46.61 33.73 26.19 21.32 - 37.38 24.36 16.55 11.79 36.74
BN-TIN-Transf.+SignBT Zhou et al. (2021a) 50.80 37.75 29.72 24.32 49.54 51.42 37.26 27.76 21.34 49.31
MMTLB Chen et al. (2022a) $\checkmark$ 53.97 41.75 33.84 28.39 52.65 53.31 40.41 30.87 23.92 53.25
TS-SLT Chen et al. (2022b) $\checkmark$ 54.90 42.43 34.46 28.95 53.48 55.44 42.59 32.87 25.79 55.72
SLTUNET Zhang et al. (2023b) $\checkmark$ 52.92 41.76 33.99 28.47 52.11 54.98 41.44 31.84 25.01 54.08
Weakly Gloss-free TSPNet Li et al. (2020b)$\ddagger$ $\checkmark$ 36.10 23.12 16.88 13.41 34.96 17.09 8.98 5.07 2.97 18.38
GASLT Yin et al. (2023) $\checkmark$ 39.07 26.74 21.86 15.74 39.86 19.90 9.94 5.98 4.07 20.35
ConSLT Fu et al. (2023) $\checkmark$ - - - 21.59 47.69 - - - - -
Gloss-free CSGCR Zhao et al. (2021) 36.71 25.40 18.86 15.18 38.85 - - - - -
GFSLT-VLP Zhou et al. (2023) $\checkmark$ 43.71 33.18 26.11 21.44 42.29 39.37 24.93 16.26 11.00 36.44
FLa-LLM Chen et al. (2024) $\checkmark$ 46.29 35.33 28.03 23.09 45.27 37.13 25.12 18.38 14.20 37.25
Sign2GPT Wong et al. (2024) $\checkmark$ 49.54 35.96 28.83 22.52 48.90 41.75 28.73 20.60 15.40 42.36
SignLLM Gong et al. (2024) $\checkmark$ 45.21 34.78 28.05 23.40 44.49 39.55 28.13 20.07 15.75 39.91
@@
Table 1: Performance comparison on the PHOENIX14T and CSL-Daily datasets. “Vis. Ft.” denotes the visually fine-tuned on sign language datasets. $\ddagger$ denotes results reproduced by Yin et al. for CSL-Daily. The best results are highlighted as bold, and the second-best are underlined.
@@ 4.1 Implementation Details
For SE and ME, we use CLIP ViT-L/14 Radford et al. (2021) and VideoMAE-L/16 Tong et al. (2022), respectively. To extract the spatial features, the sign images are interpolated to multiple scales, such as $224\times 224$ and $448\times 448$. For each scale, larger images are split into sub-images of regular size ($224\times 224$) and processed individually. These features from the sub-images are then pooled and concatenated with features from the original representation. For the motion features, each clip consists of 16 frames, based on the findings from Wilbur (2009), which suggests that this frame interval captures a single sign. We set the stride $s$ between consecutive clips to 8. We use FlanT5-XL Chung et al. (2024) as the LLM for PHOENIX14T and How2Sign, while mT0-XL Muennighoff et al. (2022) is used for the CSL-Daily. During the warm-up phase with VT-Align, we use 4K steps on PHOENIX14T and CSL-Daily, and 15K steps on How2Sign. Additional implementation details can be found in Appendix A.
4.2 Experimental Settings
@@
We evaluated our method on three sign language datasets: PHOENIX14T Camgoz et al. (2018), CSL-Daily Zhou et al. (2021a), and How2Sign Duarte et al. (2021). PHOENIX14T is a German Sign Language dataset comprising 8,257 samples and a vocabulary of 2,887 German words. CSL-Daily is a Chinese Sign Language dataset with 20,654 samples and a 2,343 Chinese characters. How2Sign focuses on American Sign Language and includes 35,191 samples with a vocabulary of 16K English words. Detailed dataset statistics are provided in Appendix B.
@@
We report BLEU via SacreBLEU Papineni et al. (2002); Post (2018)222nrefs:1|case:mixed|eff:no|tok:13a/zh|smooth:exp|version:2.2.1 and ROUGE-L Lin and Och (2004). BLEU-n assesses translation precision by evaluating n-grams. ROUGE-L measures text similarity by calculating the F1 score based on the longest common subsequences between predicted and reference texts. We also report BLEURT Sellam et al. (2020) from the BLEURT-20 checkpoint [^3] https://huggingface.co/lucadiliello/BLEURT-20, which has been shown to correlate well with human judgments.
A Taxonomy of SLT.
In Section 2, we explored gloss-free methods, including those that incorporate gloss-supervised visual encoders. Although these approaches have traditionally been categorized as gloss-free, we argue that they should more accurately be described as weakly gloss-free due to their dependence on gloss-annotated data. This classification is detailed in Table 3.6. Specifically, methods such as TSPNet Li et al. (2020b), GASLT Yin et al. (2023), ConSLT Fu et al. (2023), GloFE-VN Lin et al. (2023), and OpenSLT Tarrés et al. (2023) rely on sign features extracted by visual encoders trained on continuous or isolated sign language recognition (SLR) datasets.
@@ 4.3 Comparison with State-of-the-Art
Results on PHOENIX14T and CSL-Daily.
We first compared our method with both gloss-based and gloss-free methods on PHOENIX14T. As shown in Table 3.6, most previous methods rely on the domain-specific fine-tuning of their visual encoders. By contrast, our method demonstrates consistent improvements across all reported metrics on PHOENIX14T without such fine-tuning. The only exception is ROUGE, where we achieved the second-best result. Specifically, the improvement on BLEU-4 is by a margin of 0.92, representing a 3.93% increase over SignLLM Gong et al. (2024). On CSL-Daily, which covers a broader range of topics than PHOENIX14T, the performance gains are even more pronounced. Our method achieved a margin increase of 4.8 in BLEU-4, reflecting a 30.41% improvement over SignLLM.
@@
Next, we evaluated our method on How2Sign, which poses greater challenges than PHOENIX14T due to its broader open-domain nature, longer sign videos, and larger vocabulary. As shown in Table 3.6, our method outperformed previous methods across all reported metrics. Specifically, we achieved a 0.45 margin in BLEU-4 which represents a 4.66% improvement over Fla-LLM Chen et al. (2024). We see a performance gain in BLEURT, reaching 2.93, which is 7.46% higher than SSVP-SLT Rust et al. (2024) under the non-scaled dataset setting.
@@
Table 3: Ablation study of the main component.
Models #Trainable Params #Total Params B4
w/o LLM 60.5M 60.5M 6.35
mBART-L Liu et al. (2020) 680M 680M 10.94
mT0-XL Muennighoff et al. (2022) 23.5M 3.5B 24.23
Llama-2 Touvron et al. (2023) 32.4M 7B 13.86
@@
Table 4: Ablation study of the impact of LLM.
@@
Figure 4: Ablation study for SE and ME. $S^{2}$ represents Scaling on Scales, and $s$ denotes stride size. Note that the presented results do not include VT-Align.
@@
To assess the quality of sign representations, following Ye et al. (2023), we employed the Kernel Density Estimation (KDE) to estimate the probability density functions of embeddings from GFSLT-VLP and our method on PHOENIX14T. Note that we reproduced GFSLT-VLP using the official code [^4] https://github.com/zhoubenjia/GFSLT-VLP. As shown in Table 5, our method produced more compact and confident representations than GFSLT-VLP. More details on the KDE process are provided in Appendix A.4.
@@ 4.4 Ablation Study
To further evaluate our method, we conducted extensive ablation experiments on PHOENIX14T, the most widely used sign language dataset. Additional results can be found in Appendix C.
@@
We begin by evaluating the impact of the key components in our framework: SE, ME, and VT-Align. As shown in Table 4.3, using SE or ME individually results in lower performance, with ME performing the worst in terms of BLEU-4. However, combining SE and ME leads to an overall improvement. Notably, when VT-Align is integrated with SE, the performance rises, nearly matching Sign2GPT (22.52 vs. 22.49). The best results are achieved when all components (SE, ME, and VT-Align) are used together, yielding the highest scores across all metrics. This highlights the importance of each component in enhancing overall performance of SpaMo.
@@
Next, we explored the impact of different types of LLMs by replacing our model, as shown in Table 4.3. We compared five models, each with a different number of parameters: our method without pre-trained weights, mBART-L, mT0-XL, Flan-T5-XL, and Llama-2. Among these, Flan-T5-XL achieves the best performance, though it nearly ties with mT0-XL. Interestingly, despite its larger parameter count, Llama-2, which is also employed in SignLLM, does not outperform the others. This finding aligns with the observations of Zhang et al., suggesting that scaling up LMs does not always lead to better performance.
@@
Finally, we evaluated the effect of $S^{2}$ and the gap between neighboring clips on SE and ME, respectively. As shown in Figure 4, $S^{2}$ substantially improves translation performance, highlighting its effectiveness in helping SE capture more spatial details. Additionally, our analysis reveals that a stride size of 8 between neighboring clips yields the best results, suggesting that this stride size optimally aids ME in extracting the motion dynamics.
@@
Figure 5: The t-SNE visualization of sign features. Different colors represent features with distinct semantics, while gray points are other categories not listed.
@@ 4.5 Qualitative Analysis
Table 6 presents two example translations on PHOENIX14T, comparing our method with GFSLT-VLP, the only other publicly available baseline. In the first example (top), our method provides an accurate translation, whereas GFSLT-VLP fails to capture the correct semantic meaning. In the second example (bottom), our method again produces a precise translation, while GFSLT-VLP introduces errors, resulting in incorrect information. These examples demonstrate the superior accuracy of our method in generating reliable translations. More translation examples are shown in Appendix C.4.
@@
We performed an additional analysis to explore how the LLM interprets the sign videos. For each visual feature, we identified the word with the shortest distance in the LLM’s embedding space, representing the closest match. Further details can be found in Appendix A.5. Figure 5 shows the t-SNE visualization of each sign feature mapped to the corresponding word. We observed that certain visual features align closely with specific words, which likely represents the semantic concepts that the LLM associates with these features. In other words, these words represent the LLM’s interpretation or labeling of the visual content. We refer to these mapped words as “visual tokens”. We further compared these visual tokens with the ground-truth glosses as shown in Table 7. To ensure a clearer and more accurate semantic comparison, repetitive words were removed from the visual tokens. Surprisingly, the LLM’s interpretation of the sign videos is similar to the glosses, though not perfectly aligned. This suggests that the LLM has learned to link particular video patterns with specific textual concepts, explaining why those words cluster near the visual features in the embedding space. Additionally, we found that the visual tokens capture words that are present in the translation but not in the glosses. This finding suggests that visual tokens may provide a more comprehensive representation than current glosses, potentially broadening their scope beyond what has been traditionally documented.